OpenStudy (anonymous):

Plz help me f(x) = x2 − 2x + 9 and g(x) = 8 − x, what is (f o g)(−4)?

OpenStudy (anonymous):

g(x) 8-(-4)=12

TheSmartOne (thesmartone):

lets not keep making posts again and again >.<

TheSmartOne (thesmartone):

ok so you got the 12 part

OpenStudy (anonymous):

Yeah

TheSmartOne (thesmartone):

now lets plug in x=12 into f(x) = x^2 − 2x + 9 so f(12) = \(\sf (12)^2 -2(12) + 9\) =?

OpenStudy (anonymous):

129

OpenStudy (anonymous):

Thank U Smartone Methsucks

TheSmartOne (thesmartone):

and that's your final answer :)

TheSmartOne (thesmartone):

haha, I put Meth sucks because that rhymed with Beth sucks

OpenStudy (anonymous):

It did

TheSmartOne (thesmartone):

xD

OpenStudy (anonymous):

Thank You for helping me have a nice day

TheSmartOne (thesmartone):

you too :)